# Off-equatorial deflections and gravitational lensing. II. In general stationary and axisymmetric spacetimes

### v1

@@ Abstract
In this work, we develop a general perturbative procedure to find the off-equatorial plane deflections in the weak deflection limit in general stationary and axisymmetric spacetimes satisfying some common variable separation conditions. Deflections of both null and timelike rays with the finite distance effect of the source and detector taken into account are obtained as dual series of $M/r_{0}$ and $r_{0}/r_{s,d}$. These deflections allow a set of exact gravitational lensing equations from which the images’ apparent angular positions are solved. The method and general results are then applied to the Kerr-Newmann, Kerr-Sen, and rotating Simpson-Visser spacetimes to study the effect of the spin and characteristic (effective) charge of the spacetimes and the source altitude on the deflection angles and image apparent angles. It is found that in general, both the spacetime spin and charge only affect the deflections from the second non-trivial order, while the source altitude influences the deflection from the leading order. Because of these, it is found that in gravitational lensing in realistic situations, it is hard to measure the effects of the spacetime spin and charge from the images’ apparent locations. We also presented the off-equatorial deflections in the rotating Bardeen, Hayward, Ghosh, and Tinchev black hole spacetimes.
@@ I Introduction
The simplest scenario of signal deflection and GL is those of light rays in static and spherically symmetric (SSS) spacetime, or in the equatorial plane of stationary and axisymmetric (SAS) spacetimes in the weak deflection limit (WDL). With the fast developments of astroparticle physics [8, 9], gravitational wave detection [10] and black hole (BH) imaging [11, 12] however, enormous effort has been devoted to the extension to the deflection and GL of timelike signals [14, 13], with finite source and detector distance [15], and in the strong deflection limit [16, 17, 18]. Different analytical methods were also developed, including the perturbative methods [20, 19, 18, 21] and the more recent Gauss-Bonnet theorem-based methods [22].
However, extension to the non-equatorial deflection and GL in SAS spacetime is still rare (in SSS spacetime essentially there is no non-equatorial motion). Due to the complexity of the motion equations for the off-equatorial trajectories, only a few works have studied the deflection and GL in quasi-equatorial motion of null rays [23, 24, 25] or only numerically in the general non-equatorial case [26, 27], let alone the deflection of both null and timelike rays. In Ref. [28], the general non-equatorial deflection and GL in Kerr spacetime were studied perturbatively for the first time for both null and timelike rays in the WDL, with finite distance effect taken into account. One of the motivations of the current work is to study the condition on the form of SAS spacetime for the perturbative method to be feasible for non-equatorial deflection and GL. We will show that for many SAS spacetimes satisfying a separation condition, the perturbative method is always valid for both null and timelike rays with the finite distance effect automatically taken into account. The second motivation is to reveal the effects of various spacetime parameters on such deflections and GL. Spacetimes will be considered include the Kerr-Newmann (KN), Kerr-Sen (KS) [29], rotating Simpson–Visser (RSV) [30] and a few other ones [32, 33, 34, 31].
The paper is organized as follows. In Sec. II.1, we first introduce the basic setup of the problem, establish the equations of motion, and study a condition for the perturbative method to work. In Sec. III, we explore a perturbative method to solve the deflection angle in both $\theta$ and $\phi$ direction. The off-equatorial GL equations are then solved in Sec. IV to obtain the apparent angles of the lensed images. The method and deflection and GL results are then applied to the KN, KS, and RSV spacetimes in Sec. V, and the effects of their characteristic parameters are studied. We conclude the paper with a summary and discussion in Sec. VI. Throughout the paper we use the natural unit $G=c=1$ and the spacetime signature (,+,+,+-,+,+,+- , + , + , +).
@@ II General framework
In this section, we will derive the deflection angles in the $\theta$ and $\phi$ directions for SAS spacetimes. We will show that this is always possible when the metric functions satisfy certain conditions such that a proper separation of variables in the equations of motion, or equivalently the existence of a generalized Carter constant (GCC), can be accomplished.
@@ II.1 Preliminaries
where $t,\,r,\,\theta,\,\phi$ are the Boyer-Lindquist coordinates and $A,\,B,\,C,\,D,\,F$ are functions of $r$ and $\theta$ only. This metric allows two commutive Killing vectors
@@
where the spacelike $\psi^{\mu}$ corresponding to the rotation symmetry and timelike $\xi^{\mu}$ to the time translation. These Killing vectors generate two conserved quantities of the motion
@@
From these equations, one can obtain two first derivatives
@@
where $\kappa=0,\,-1$ for null and timelike particles respectively. Here and henceforth any function with an $(r)$ or $(\theta)$ superscript is a function of $r$ or $\theta$ only. The Hamilton-Jacobi equation is given by
@@
We then seek metrics that allow this equation, after multiplied by some proper total factor function $G(r,\theta)$, to be separable into $r$ and $\theta$ dependent parts. This can be accomplished if the metric functions $A,\,B,\,C,\,D,\,F$, and factor $G$ can cast the left-hand sides of the following equations into their right-hand sides
@@
where $X\in\{A,\,B,\,C\}$. Note here the condition (10c) is for the separability of the $\kappa=-1$ case and unnecessary for null signals. Condition (10a) implies
@@
In practice, the functions on the right-hand sides of Eq. (10) as well as $G(r,\theta)$ can be simply read off from the left-hand sides and Eq. (11). One might also notice that there is a freedom of a multiplicative constant in functions $\mathcal{D}(r)$ and $\mathcal{F}(\theta)$, and additive constant freedom in each pair of functions $X^{(r)}$ and $X^{(\theta)}$. Indeed one can show that these freedoms will be canceled out in the final equations of motion (17) and (18) and therefore not affect the physics. Let us also point out that many SAS spacetimes, including the Kerr spacetime, satisfy these conditions (10).
@@
where we have defined two compact functions $R(r)$ and $\Theta(\theta)$ to simplify the notation. We remind that once the metric functions are known, these two functions are determined too. The function $S^{(r)}$ and $S^{(\theta)}$ therefore can be solved and the action (7) becomes
@@ II.2 Defining the deflection $\Delta\phi$ and $\Delta\theta$
Before we dive further into the technical details however, let us also remind that since we are in a WDL, it is reasonable to assume that during the propagation of the signal, it only experiences one periapsis with radius $r_{0}\gg M$ where $M$ is the characteristic length scale of the spacetime, and one extreme value of the azimuth angle $\theta_{m}$. I.e.,
@@
and $s_{2}=\pm 1$ is introduced when solving a quadratic equation. These relations connect the motion constants $(E,\,L,\,K)$ with $(r_{0},\,\theta_{m})$. In Sec. III, we will use $(r_{0},\,\theta_{m})$ to tread for $(L,\,K)$ since the latter are less intuitive and usually harder to measure in astronomy.
@@
This relation allows us to solve $\theta_{d}$ once $\theta_{s}$, the spacetime, and other kinetic variables are known. Therefore from this, we will be able to find the deflection in the $\theta$ direction as defined in Eq. (19).
@@ III.1 The Perturbative Method
The main idea of the perturbative method is selecting appropriate expansion parameter(s) and expanding the interands into simpler series so that the integrations become doable. In the WDL, there is a natural small parameter $1/r_{0}$ suitable for this purpose. When expanding the integrands in Eqs. (28) and (29), one can also anticipate that the expansion coefficients will explicitly depend on the asymptotic behavior of the metric functions. After a short survey of the applicable spacetime metrics of our method, we found the following expansions can be assumed for the functions $X^{(\mu)}~{}(\mu=r,\,\theta)$ and $\mathcal{D}(r)$ and $\mathcal{F}(\theta)$
@@
Some higher-order terms are present in Appendix A. For the integrals over $p$ in Eqs. (33) and (34), we can show that their integrands are always of the form $\text{polynomrial}(p)/\left(1-p^{2}\right)^{n+1/2}~{}(n=0,\,1,\,\cdots)$ and therefore integrable [28]. For the integral over $c$ in these equations, their integrands are always of the form $\text{polynomrial}(c)/\left[(1-c^{2})^{n}\sqrt{c_{m}^{2}-c^{2}}\right]~{}(n=0,\,1)$ and therefore also integrable.
@@ III.2 Deflection angles
Eqs. (41) and (44) are two important results of the work and a few comments are in order for them. Firstly, these results apply to general SAS spacetimes that allow the existence of a GCC. This includes many well-known spacetimes such as the Kerr, KN, and all SSS ones, which can be obtained by setting all $b_{n}=0$ for $n\geq 2$. Secondly, they work for both light rays (setting $\kappa=0$) and timelike particles (setting $\kappa=-1$). Indeed we can show that the $E\to\infty$ limits of these results for timlike signals equal exactly their values for null rays. Thirdly, these deflections also take into account the finite distance effect of the source and detector. This effect could be important when studying the GL effect. Setting $p_{s,d}$ to zero, the infinite distance version of the deflections can be obtained. Fourthly, these deflections work for both non-equatorial and equatorial trajectories. For the former, setting $\theta_{s,d}\to\pi/2$, we have verified that the deflection angle $\Delta\phi$ reduces to its value on the equatorial plane in the Kerr spacetime [19]. For the latter case, these formulas do not rely on any near-equatorial plane approximation, i.e., $\theta_{s}$ and $\theta_{d}$ can be far from $\pi/2$. Last but not least, the deflections (41) and (44) can be further expanded around small $p_{s,d}$ if the source and detector are far away, i.e., $r_{s,d}\gg r_{0}$, and a dual series form will be obtained. Such a form will be more appealing from the application point of view and we will only do this in Sec. V.
@@ IV Gravitational lensing
Since we now have the deflection angles for arbitrary inclination angles of the signals, we can study the GL in such SAS spacetime in the off-equatorial plane too.
Then since our deflection angles (41) and (44) contain the finite distance effect, we can naturally establish the following GL equations
@@
It is found that in general, there will be two sets of $(r_{0},\,\theta_{m})$ that allow the signal to reach the detector. Most of the time when the deflection $(\delta\theta,\,\delta\phi)$ is not small, these two solutions will have opposite orbital rotation direction $s_{2}$ and the $s_{1}=\mathrm{sign}(\theta_{m})$ equals $s_{2}$ for each solution. Only when the source, lens and detector are very aligned (deflection less than $10^{-6\prime\prime}$ in the Sgr A* situation considered in Sec. V.1) and the spacetime spin is large, there could exist exceptions (see also Ref. [28]). Therefore we will label these two solutions as $(r_{0+},\,\theta_{m+})$ and $(r_{0-},\,\theta_{m-})$ to represent the prograde and retrograde rotating signals respectively.
@@ V.2 Deflections and GL in KS spacetime
KS BH is a kind of rotating and charged black hole in the four-dimensional heterotic string theory [29]. Although both the strong [45] and weak [46] deflection limits of GL effects have been studied in this spacetime using approaches different from ours, these studies are either in the (quasi-)equatorial plane or did not express the deflections in terms of the original source and kinetic variables. Here we will extend them to the general non-equatorial case, with finite distance effect and timeline effect taken into account.
@@ V.3 Deflections and GL in RSV spacetime
where $\Theta_{\pm}$ is still given by Eq. (55) while its $(L_{\pm},\,K_{\pm})$ depends on $(r_{0\pm},\,\theta_{m\pm})$ through Eqs. (22) and (23). The approximation of these apparent angles are
@@
In Fig. 5 (b), we plotted the dependences of the deflections $(\Delta\phi_{\mathrm{R}},\,\Delta\theta_{\mathrm{R}})$ and the apparent angles $(\alpha_{\pm\mathrm{R}},\,\beta_{\pm\mathrm{R}})$ on $\hat{l}$. Similar to the effect of $\hat{Q}$ in KN spacetime, here $\hat{l}$ also decreases $r_{0\pm}$ and consequently the apparent angles of the images. The amount of reduction of either $\alpha_{\pm R}$ or $\beta_{\pm R}$ as $\hat{l}$ grows to 2.5 is comparable to that for $\hat{Q}$ in the KN case or $\hat{b}$ in the KS case.

### v3

@@ Abstract
In this work, we develop a general perturbative procedure to find the off-equatorial plane deflections in the weak deflection limit in general stationary and axisymmetric spacetimes, allowing the existence of the generalized Carter constant. Deflections of both null and timelike rays, with the finite distance effect of the source and detector taken into account, are obtained as dual series of $M/r_{0}$ and $r_{0}/r_{s,d}$. These deflections allow a set of exact gravitational lensing equations from which the images’ apparent angular positions are solved. The method and general results are then applied to the Kerr-Newmann, Kerr-Sen, and rotating Simpson-Visser spacetimes to study the effect of the spin and characteristic (effective) charge of the spacetimes and the source altitude on the deflection angles and image apparent angles. It is found that, in general, both the spacetime spin and charge only affect the deflections from the second non-trivial order, while the source altitude influences the deflection from the leading order. Because of this, it is found that, in gravitational lensing in realistic situations, it is hard to measure the effects of the spacetime spin and charge from the images’ apparent locations. We also presented the off-equatorial deflections in the rotating Bardeen, Hayward, Ghosh, and Tinchev black hole spacetimes.
@@ I Introduction
The simplest scenario of signal deflection and GL is that of light rays in static and spherically symmetric (SSS) spacetimes, or in the equatorial plane of stationary and axisymmetric (SAS) spacetimes in the weak deflection limit (WDL). With the fast developments of astroparticle physics [8, 9], gravitational wave detection [10], and black hole (BH) imaging [11, 12], enormous effort has been devoted to the extension of the deflection and GL of timelike signals [14, 13, 15, 16, 17, 18], with finite source and detector distance [19], and in the strong deflection limit [20, 21, 22]. Different analytical methods were also developed, including the perturbative methods [24, 23, 22, 25] and the more recent Gauss-Bonnet theorem-based methods [26, 16, 17, 18].
However, extension to the non-equatorial deflection and GL in SAS spacetime is still rare (in SSS spacetime essentially there is no non-equatorial motion), except in Kerr [27, 28, 29, 30, 31, 32, 33] and Kerr-Newmann (KN) [34, 35] spacetime. Due to the complexity of the motion equations for the off-equatorial trajectories, only a few works have studied the deflection and GL in quasi-equatorial motion of null rays [36, 37] or only numerically in the general non-equatorial case [38, 39] in other spacetimes, let alone the deflection of both null and timelike rays. In Ref. [33], the general non-equatorial deflection and GL in Kerr spacetime were studied perturbatively for the first time for both null and timelike rays in the WDL, with finite distance effect taken into account. One of the motivations of the current work is to study the condition under the form of SAS spacetime for the perturbative method to be feasible for non-equatorial deflection and GL. We will show that for many SAS spacetimes satisfying a separation condition that allows a generalized Carter constant (GCC), the perturbative method is always valid for both null and timelike rays with the finite distance effect automatically taken into account. The second motivation is to reveal the effects of various spacetime parameters on such deflections and GL. Spacetimes will be considered include the KN, Kerr-Sen (KS) [40], rotating Simpson–Visser (RSV) [41] and a few other ones [43, 44, 45, 42].
The paper is organized as follows. In Sec. II.1, we first introduce the basic setup of the problem, establish the equations of motion, and study a condition for the perturbative method to work. In Sec. III, we explore a perturbative method to solve the deflection angle in both $\theta$ and $\phi$ direction. The off-equatorial GL equations are then solved in Sec. IV to obtain the apparent angles of the lensed images. The method and deflection and GL results are then applied to the KN, KS, and RSV spacetimes in Sec. V, and the effects of their characteristic parameters are studied. We conclude the paper with a summary and discussion in Sec. VI. Throughout the paper, we use the natural unit $G=c=1$ and the spacetime signature (,+,+,+-,+,+,+- , + , + , +).
@@ II General framework
In this section, we will derive the deflection angles in the $\theta$ and $\phi$ directions for SAS spacetimes. We will show that this is always possible when the metric functions satisfy certain conditions, such that a proper separation of variables in the equations of motion, or equivalently the existence of a GCC, can be accomplished.
@@ II.1 Preliminaries
where $t,\,r,\,\theta,\,\phi$ are the Boyer-Lindquist coordinates and $A,\,B,\,C,\,D,\,F$ are functions of $r$ and $\theta$ only. This metric allows two commutative Killing vectors
@@
where the spacelike $\psi^{\mu}$ corresponding to the rotation symmetry and timelike $\xi^{\mu}$ to the time translation. These Killing vectors correspond to two conserved quantities of the motion
@@
The asymptotic velocity $v$ here stands for the magnitude of the spatial component of the four-velocity of the test particle. From these equations, one can obtain two first derivatives
@@
where $\kappa=0,\,-1$ for null and timelike particles respectively. Here and hereafter, any function with an $(r)$ or $(\theta)$ superscript is a function of $r$ or $\theta$ only. The Hamilton-Jacobi equation is given by
@@
We then seek metrics that allow this equation, after multiplied by some proper total factor function $G(r,\theta)$, to be separable into $r$ and $\theta$ dependent parts. This can be accomplished if the metric functions $A,\,B,\,C,\,D,\,F$, and factor $G$ can cast the left-hand sides of the following equations into their right-hand sides [46]
@@
where $X\in\{A,\,B,\,C\}$. Note here that the condition (10c) is for the separability of the $\kappa=-1$ case and unnecessary for null signals. Condition (10a) implies
@@
In practice, the functions on the right-hand sides of Eq. (10) as well as $G(r,\theta)$ can be read off from the left-hand sides and Eq. (11). One might also notice that there is a freedom of a multiplicative constant in functions $\mathcal{D}(r)$ and $\mathcal{F}(\theta)$, and additive constant freedom in each pair of functions $X^{(r)}$ and $X^{(\theta)}$. Indeed one can show that these freedoms will be canceled out in the final equations of motion (17) and (18) and therefore not affect the physics. Let us also point out that many SAS spacetimes, including the Kerr spacetime, satisfy these conditions (10).
A few comments about the variable separation condition (10) might be useful for their clear understanding here. What will be shown in this work is that the spacetimes satisfying the condition (10) can always be treated using our method, while those spacetimes not satisfying (10) are not treatable by the method developed in this work. In this sense, condition (10) is both a sufficient and necessary condition for the applicability of our method. However, Eq. (10) is only a sufficient condition for the separability of the equations of motion and we are not able to prove that it is also a necessary condition, although we can not provide any counter-example either. That is to say, it is unclear to us whether there exist spacetimes (unknown to us) that do not satisfy condition (10) but still allow the separation of its variables.
@@
where we have defined two compact functions $R(r)$ and $\Theta(\theta)$ to simplify the notation. Once the metric functions are known, these two functions are determined too. The function $S^{(r)}$ and $S^{(\theta)}$ therefore can be solved and the action (7) becomes
@@ II.2 Defining the deflection $\Delta\phi$ and $\Delta\theta$
In the WDL, it is reasonable to assume that during the propagation of the signal, it only experiences one periapsis with radius $r_{0}\gg M$, where $M$ is the characteristic length scale of the spacetime, and one extreme value of the azimuth angle $\theta_{m}$. I.e.,
@@
and $s_{2}=\pm 1$ is introduced when solving a quadratic equation. These relations connect the motion constants $(E,\,L,\,K)$ with $(r_{0},\,\theta_{m})$. In Sec. III, we will use $(r_{0},\,\theta_{m})$ to replace $(L,\,K)$ since the latter are less intuitive and usually harder to measure in astronomy. For example, $r_{0}$ for the bending by the Sun can be approximated by the solar radius.
@@
This relation allows us to solve $\theta_{d}$ once $\theta_{s}$, the spacetime, and other kinetic variables are known. Therefore from this, we can find the deflection in the $\theta$ direction as defined in Eq. (19).
@@ III.1 The Perturbative Method
The main idea of the perturbative method is selecting appropriate expansion parameter(s) and expanding the integrands into simpler series so that the integrations become doable. In the WDL, there is a naturally small parameter $1/r_{0}$ suitable for this purpose. When expanding the integrands in Eqs. (28) and (29), one can also anticipate that the expansion coefficients will explicitly depend on the asymptotic behavior of the metric functions. After a short survey of the applicable spacetime metrics of our method, we found the following expansions can be assumed for the functions $X^{(\mu)}~{}(\mu=r,\,\theta)$ and $\mathcal{D}(r)$ and $\mathcal{F}(\theta)$
@@
Some higher-order terms are presented in Appendix A. For the integrals over $p$ in Eqs. (33) and (34), we can show that their integrands are always of the form $\text{polynomrial}(p)/\left(1-p^{2}\right)^{n+1/2}~{}(n=0,\,1,\,\cdots)$ and therefore integrable [33]. For the integral over $c$ in these equations, their integrands are always of the form $\text{polynomrial}(c)/\left[(1-c^{2})^{n}\sqrt{c_{m}^{2}-c^{2}}\right]~{}(n=0,\,1)$ and therefore also integrable.
@@ III.2 Deflection angles
Eqs. (41) and (44) are two important results of the work and a few comments are for them. Firstly, these results apply to general SAS spacetimes that allow the existence of a GCC. This includes many well-known spacetimes such as the Kerr, KN, and all SSS ones, which can be obtained by setting all $b_{n}=0$ for $n\geq 2$. Secondly, they work for both light rays (setting $\kappa=0$) and timelike particles (setting $\kappa=-1$). Indeed we can show that the $E\to\infty$ limits of these results for timlike signals equal exactly their values for null rays. Thirdly, these deflections also take into account the finite distance effect of the source and detector. This effect could be important when studying the GL effect. Setting $p_{s,d}$ to zero, the infinite distance version of the deflections can be obtained. Fourthly, these deflections work for both non-equatorial and equatorial trajectories. For the former, setting $\theta_{s,d}\to\pi/2$, we have verified that the deflection angle $\Delta\phi$ reduces to its value on the equatorial plane in the Kerr spacetime [23]. For the latter case, these formulas do not rely on any near-equatorial plane approximation, i.e., $\theta_{s}$ and $\theta_{d}$ can be far from $\pi/2$. Last but not least, the deflections (41) and (44) can be further expanded around small $p_{s,d}$ if the source and detector are far away, i.e., $r_{s,d}\gg r_{0}$, and a dual series form will be obtained. Such a form will be more appealing from the application point of view and we will only do this in Sec. V.
@@ IV Gravitational lensing
As we’ve got the deflection angles for arbitrary inclination angles of the signals, we can study the GL in such SAS spacetime in the off-equatorial plane. Since our deflection angles (41) and (44) contain the finite distance effect, we can naturally establish the following GL equations
@@
It is found that in general, there will be two sets of $(r_{0},\,\theta_{m})$ that allow the signal to reach the detector. Most of the time when the deflection $(\delta\theta,\,\delta\phi)$ is not small, these two solutions will have opposite orbital rotation direction $s_{2}$ and the $s_{1}=\mathrm{sign}(\cos(\theta_{m}))$ equals $s_{2}$ for each solution. Only when the source, lens and detector are very aligned (deflection less than $10^{-6\prime\prime}$ in the Sgr A* situation considered in Sec. V.1) and the spacetime spin is large, there could exist exceptions (see also Ref. [33]). Therefore we will label these two solutions as $(r_{0+},\,\theta_{m+})$ and $(r_{0-},\,\theta_{m-})$ to represent the prograde and retrograde rotating signals respectively. In this paper, by prograde and retrograde, we always and only mean that the trajectories are rotating anticlockwise and clockwise around the $+\hat{z}$ directions respectively. No retrolensing is involved because we only discuss the weak deflection cases in this work.
@@ V.2 Deflections and GL in KS spacetime
KS BH is a kind of rotating and charged black hole in the four-dimensional heterotic string theory [40]. Although both the strong [56] and weak [36] deflection limits of GL effects have been studied in this spacetime using approaches different from ours, these studies are either in the (quasi-)equatorial plane or did not express the deflections in terms of the original source and kinetic variables. Here we will extend them to the general non-equatorial case, with finite distance effect and timelike effect taken into account.
@@ V.3 Deflections and GL in RSV spacetime
where $\Theta_{\pm}$ is still given by Eq. (55) while its $(L_{\pm},\,K_{\pm})$ depends on $(r_{0\pm},\,\theta_{m\pm})$ through Eqs. (22) and (23). The approximations of these apparent angles are
@@
In Fig. 5 (b), we plotted the dependences of the deflections $(\Delta\phi_{\mathrm{R}},\,\Delta\theta_{\mathrm{R}})$ and the apparent angles $(\alpha_{\pm\mathrm{R}},\,\beta_{\pm\mathrm{R}})$ on $\hat{l}$. Unlike the effect of $\hat{Q}$ in KN spacetime, here $\hat{l}$ increases $r_{0\pm}$ and consequently the apparent angles of the images. The amount of reduction of either $\alpha_{\pm R}$ or $\beta_{\pm R}$ as $\hat{l}$ grows to 2.5 is comparable to that for $\hat{Q}$ in the KN case or $\hat{b}$ in the KS case.